OpenStudy (anonymous):

Help! "4 times a number,subtracted fom 10, is twice the number. Find the number." I had 5(x-6) =3x x is 5 but its wrong.

OpenStudy (anonymous):

okay so the problem is 4x-10=2x right?

OpenStudy (anonymous):

subtracted from 10 - 4 times the number = 2 times the number 10 - 4x = 2x 10 = 6x x = 5/3

OpenStudy (anonymous):

thats what i had, but said i was wrong... brc39:: that makes sense, but why 10-4x not 4x-10

OpenStudy (anonymous):

thats what i was thinking. but if it is the way i said then it would be x=5

OpenStudy (anonymous):

typo meant 4

OpenStudy (anonymous):

It say 4x is subtracted FROM 10, so it has to be 10-4x. If it said 4 times the number subtracting 10 or minus 10 it would 4x-10. Be careful for little words like from and such to give order.

OpenStudy (anonymous):

4x-10=2x -4x -4x -10= -2x -10/2x= -2x/-2x 5=x